OpenStudy (anonymous):

You are designing a new container for a powdered laundry detergent. You are considering the cylindrical container with a diameter of 14 inches and a height of 18 inches. Find the volume of this container to the nearest cubic inch.

OpenStudy (anonymous):

\[ V = \pi r^2 h = 3.14 \times 14^2 \times 18 = 11077.92 \approx 11078 \]